OpenStudy (anonymous):

Just another cute integral: \[\int |x|^3 \; dx \] and yes this is indefinite ;)

OpenStudy (anonymous):

Hmm \[ \int |x|^2 = \left\{\begin{array}{kcc}\frac{1}{4}\cdot x^4, \quad x>0 \\ 0, \quad x = 0 \\ \frac{-1}{4}x^4, \quad x<0 \end{array}\right.\]Maybe?

OpenStudy (anonymous):

I have never seen a indefinite integral represented like this.

OpenStudy (anonymous):

Me neither but I never had to integrate modulus functions without limits before. It has to be like this, I am quite sure.

OpenStudy (anonymous):

\[\int\limits_{}^{}|x|^3 dx = \frac{x^3*|x|}{4}+c\] I hope this is right @FoolForMath and congrats on being a new moderator

OpenStudy (anonymous):

This can't be right, |x| stays positive but the x^3 can get negative.

OpenStudy (anonymous):

I am not sure though. and I can be wrong.

OpenStudy (anonymous):

For negative x , |x| = -x and therfore integral value would be positive. Though I can be wrong @ this

OpenStudy (anonymous):

For negative x, x^3 * |x| = (-x)^3 (-(-x)) = -x^4

OpenStudy (anonymous):

What's wrong piece wise representation?

OpenStudy (anonymous):

wait... my piece wise is gonna give same result as yours

OpenStudy (anonymous):

so, either we are both wrong or right

OpenStudy (anonymous):

\[\huge \int\limits\limits_{}^{}|x|^3 dx = {sgn(x)}* \frac{x^4}{4}+c\]

OpenStudy (anonymous):

and your integral does seems better

OpenStudy (anonymous):

ohh yeah

OpenStudy (anonymous):

sgn(x) how could i not think about it!?!?! :(

OpenStudy (anonymous):

so, there are multiple ways of representing same integral?

OpenStudy (anonymous):

because every integral on this page is gonna get you the same result

OpenStudy (anonymous):

so, how do you rule out others? and make one right.

OpenStudy (anonymous):

@Ishaan94 , Is my integral right because substitute -x and check please

OpenStudy (anonymous):

-x^4/4 is what we will get. sgn(-x)x^4/4 = -1* x^4/4

OpenStudy (anonymous):

Yes and why is it negative??

OpenStudy (anonymous):

I have confused myself lolol

OpenStudy (anonymous):

i am confused myself and blabbering :/

OpenStudy (anonymous):

Same here LOL. However I wolframed it and it is giving my equation. The sgn(x) one. Ditto . ??

OpenStudy (anonymous):

Ok I got it. Taking example always helps

OpenStudy (anonymous):

In that case how is my first equation wrong , the very first one I mentioned ??

OpenStudy (anonymous):

\[\int_{-1}^1 |x|^3 = \int_{-1}^0 -x^3 + \int_{0}^1 x^3 = -(-1/4) + \frac14 = \frac12 \] \[\left[sgn(x) \cdot \frac{x^4}{4}\right]_{-1}^1 = \frac{1}4 + \frac14\]

OpenStudy (anonymous):

Your first integral was right as well

OpenStudy (anonymous):

My piece wise representation is pelletty :/

OpenStudy (anonymous):

Haha. Maths always confuses us even when we have the right answers . :P

OpenStudy (anonymous):

yeah hahaha i never had to integrate such indefinite integrals before.

OpenStudy (anonymous):

Why isn't FoolForMath replying?

OpenStudy (anonymous):

Only 2 people know the answer to your question @Ishaan94 , @FoolForMath and god himself :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Lol, I had the answer but I only figured it out why is it working. Thanks to M.se again ;)

OpenStudy (anonymous):

@God

OpenStudy (anonymous):

0 fans lol

OpenStudy (anonymous):

I beat God! by 455 fans

OpenStudy (anonymous):

LOLOL

OpenStudy (anonymous):

So, the answer is \[ \frac 1 4 |x|^3 \cdot x +C \]

OpenStudy (anonymous):

Does it matter, @FoolForMath the position of modulus on either x^3 or x :P